# CLMFormer: Mitigating Data Redundancy to Revitalize Transformer-based Long-Term Time Series Forecasting System

### v4

@@
Abstract
Long-term time-series forecasting (LTSF) plays a crucial role in various practical applications. Transformer and its variants have become the de facto backbone for LTSF, offering exceptional capabilities in processing long sequence data. However, existing Transformer-based models, such as Fedformer and Informer, often achieve their best performances on validation sets after just a few epochs, indicating potential underutilization of the Transformer’s capacity. One of the reasons that contribute to this overfitting is data redundancy arising from the rolling forecasting settings in the data augmentation process, particularly evident in longer sequences with highly similar adjacent data. In this paper, we propose a novel approach to address this issue by employing curriculum learning and introducing a memory-driven decoder. Specifically, we progressively introduce Bernoulli noise to the training samples, which effectively breaks the high similarity between adjacent data points. To further enhance forecasting accuracy, we introduce a memory-driven decoder. This component enables the model to capture seasonal tendencies and dependencies in the time-series data and leverages temporal relationships to facilitate the forecasting process. The experimental results on six real-life LTSF benchmarks demonstrate that our approach can be seamlessly plugged into varying Transformer-based models, with our approach enhancing the LTSF performances of various Transformer-based models by maximally 30%.
Index Terms:
@@ Abstract
I Introduction
@@
Long-term time series forecasting (LTSF) stands as a cornerstone within a multitude of practical applications for complex real-life scenarios, such as finance [1], traffic, weather [2] and hospital admission rates prediction [3], wielding the power to shape crucial outcomes. Witness the success of Transformer [4] and further applications in numerous fields [5, 6, 7, 8, 9, 10, 11], researchers have begun to utilize the Transformer and its evolutionary descendants [12, 13, 14, 15] as the de facto backbone in this intricate landscape. It showcases unparalleled prowess in handling extensive sequence data, elucidates patterns, and unveils latent structures that underscore sequential phenomena. More details, these Transformer-based LTSF systems have demonstrated strong performance, for which they can predict hundreds of time steps ahead, in contrast to other sequence processing networks and traditional methods that were mostly only capable of making predictions for 48-time points or less [16, 17, 18]. The spotlight on the Transformer architecture within LTSF illuminates a sophisticated symphony of computation, where self-attention mechanisms and hierarchical processing harmoniously intertwine. This orchestration empowers these models to decipher intricate temporal intricacies and discern long-range dependencies embedded within extensive sequences [19].
Figure 1: Illustration of the rolling forecasting setting with stride size=1.
@@
Strikingly, when playing with these Transformer-based LTSF systems [14, 19, 15, 4], we observed that these models frequently attain their apex performance on validation datasets within a remarkably abbreviated span of training epochs, sometimes even the first epoch [15], and then the performances dropped gradually. This observation hints at the latent underexploitation of the Transformer-based systems’ innate potential. The emergence of such a phenomenon can typically be attributed to a confluence of two primary factors: insufficient training data and the constrained parameter within the network architecture. However, the inherent modularity of these architectures allows for facile parameter augmentation by virtue of stacking additional Transformer modules. This architectural extensibility, coupled with the recognition that the constrained parameter space might not be the predominant bottleneck, guides our inference toward the more probable origin: the paucity of comprehensive and high-quality training data.
@@
Delving further into this phenomenon, we found that there is severe data redundancy within the training samples, stemming from the common practice of rolling forecasting settings (see Figure. 1 for more details) during data augmentation. Under the rolling forecasting setting, if the stride size is 1, then the previous and current training samples are only one data point apart from each other. Since making longer predictions typically requires longer input sequences, training samples for LTSF models have a higher degree of similarity. For example, when the model’s input length is 2, the similarity between two consecutive training inputs is 50% (1/2); however, if we increase the model’s input length to 168, the similarity between two consecutive training inputs rises to 99.4% (167/168). Therefore, the LTSF models are more likely to suffer from overfitting due to such data redundancy. One could argue that we could decrease the training sample similarities by increasing the stride size; however, the number of training samples decreases by the factor of the stride size, as the training sample numbers are calculated by:
$N_{t}=\frac{N_{d}-S_{w}+1}{S_{s}}\rm{.}$ (1)
@@
To address the aforementioned data redundancy issue, we propose CLMFormer, a novel pipeline with curriculum learning and a memory-driven decoder, designed to revitalize varying Transformer-based LTSF systems. Since data redundancy directly results from the heightened similarity present within the training data, our priority is to break this similarity while concurrently preserving the necessary training sample volume. To achieve this, we leverage the concept of curriculum learning[20, 21, 22], gradually increasing training difficulty and data variety by progressively introducing Bernoulli noise to the training data through a dynamic dropout scheme. Such a scheme reduces the propensity for overfitting while simultaneously enhancing the model’s capacity to capture sequential patterns. To further enhance the Transformer’s ability to recognize intricate sequential patterns from highly similar training data, we propose a memory-driven decoder to facilitate the forecasting process. This is accomplished by adopting a seasonal memory and memory-conditioned layer normalization operation, both of which effectively utilize past time-series information and introduce additional computations to enhance the model’s capability for complex pattern recognition. The main contributions of this paper are summarized as follows:
@@
Figure 2: An overview of the model architecture. $M_{t-1}$ represents the memory matrix from the previous prediction, $M_{t}$ represents the updated memory matrix during the current sequence prediction, which in turn is recycled to be the $M_{t-1}$ for the next prediction.
II Related Work
II-A Time Series Forecasting
@@
Time series forecasting (TSF) [23] is a critical task in various domains, encompassing domains such as finance, energy, healthcare, and more. Early approaches to TSF often revolved around traditional statistical and machine learning methods. These methods include ARIMA models [24], exponential smoothing methods [25], and seasonal decomposition of time series algorithms [26]. These techniques have been widely used due to their simplicity and interpretability. However, they often struggle to capture intricate temporal patterns and dependencies present in real-world data.
In recent years, deep learning techniques have gained traction for TSF. RNNs, including LSTM[16], GRU[27] and DeepAR[28], have demonstrated their capability to capture temporal patterns. However, these models easily suffer from error accumulation, a phenomenon where small errors generated during earlier time steps can propagate and magnify as the sequence progresses[29]. This issue can significantly impact the overall accuracy and stability of the model’s predictions, especially in long sequences or when dealing with complex relationships within the data.
II-B Transformer-based LTSF Systems
Due to the superior performances in processing sequence data, Transformer-based models have achieved state-of-the-art results in many time-series tasks [30], especially in LTSF problems [12, 19, 15, 31]. There are several areas of focus among the latter studies. Some have focused on reducing the computational costs to enable longer predictions by sparse attention, whereby the memory cost is reduced from O($L^{2}$) to as low as O($LlogL$) [32, 15], where $L$ refers to the sequence length in attention layers. Other studies aim at improving the architecture or patch representations to further enhance the Transformer’s ability to learn seasonal and intricate temporal patterns. The Autoformer model[19] introduces an innovative Auto-Correlation mechanism that enables sub-series level aggregation, effectively addressing the inherent periodicity within time series data. FEDformer[14], on the other hand, focuses on learning a sparse temporal representation through the utilization of frequency-enhanced blocks/attention. This approach involves leveraging a random subset of frequency components along with learnable complex-number parameters. Another recent development, InParformer[12], incorporates evolutionary seasonal-trend decomposition modules and an interactive parallel attention mechanism with frequency-aware attention and time-aware attention, enhancing its ability to capture meaningful patterns within time series data. PatchTST[13] adopts subseries-level patches and channel independence to improve the capability of representation learning. Despite the significant progress made, Zeng et al.[33] argued that Transformer-based LTSF models are underutilized. They proposed a simple linear model that achieved superior performance. In this paper, we revisit these models and uncover the issue attributed to data redundancy stemming from rolling forecasting settings.
III Approach
@@
In this section, we present our solutions to revitalize Transformer-based systems that specifically target the challenges for LTSF. Our approach is based upon the Transformer architecture; we aim to enhance Transformer models on LTSF problems by making improvements to mitigate data redundancy from both training optimization prospects and the architecture. The overview of our approach is illustrated in Figure. 2. Our model comprises three major elements, namely the input embedding, the main Transformer-based model, and the seasonal memory component, whereby we propose to train the model based on a progressive training strategy. We first introduce the background of employing Transformer models for LTSF and then describe the training strategy. The seasonal memory component is described in the remaining subsections.
@@
Typical Transformer-based LTSF models normally train under rolling forecasting settings with a fixed window and a stride size = 1, as shown in Figure 1. The input data at time t can be written as $X^{t}=\{x_{1}^{t},x_{2}^{t}...x^{t}_{\emph{L}_{S}}|x_{k}^{t}\in\rm{R}^{\emph{d}_{\emph{f}}}\}$, where ${L}_{S}$ is the input length and the feature dimension $d_{f}>1$. Moreover, the model output is $Y^{t}=\{y_{1}^{t},y_{2}^{t}...y^{t}_{\emph{L}_{p}}\}$, where ${L}_{p}$ is the prediction length, and the output feature dimension is flexible and is not limited to single sequence predictions.
III-A Base Model
III-A1 Input Embedding

### v5

# Mitigating Data Redundancy to Revitalize Transformer-based Long-Term Time Series Forecasting System

@@
Abstract.
Long-term time-series forecasting (LTSF) is fundamental to various real-world applications, where Transformer-based models have become the dominant framework due to their ability to capture long-range dependencies. However, these models often experience overfitting due to data redundancy in rolling forecasting settings, limiting their generalization ability particularly evident in longer sequences with highly similar adjacent data. In this work, we introduce CLMFormer, a novel framework that mitigates redundancy through curriculum learning and a memory-driven decoder. Specifically, we progressively introduce Bernoulli noise to the training samples, which effectively breaks the high similarity between adjacent data points. This curriculum-driven noise introduction aids the memory-driven decoder by supplying more diverse and representative training data, enhancing the decoder’s ability to model seasonal tendencies and dependencies in the time-series data. To further enhance forecasting accuracy, we introduce a memory-driven decoder. This component enables the model to capture seasonal tendencies and dependencies in the time-series data and leverages temporal relationships to facilitate the forecasting process. Extensive experiments on six real-world LTSF benchmarks show that CLMFormer consistently improves Transformer-based models by up to 30%, demonstrating its effectiveness in long-horizon forecasting.
@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learning algorithmsccs: Computing methodologies Modeling methodologies
1. Introduction
@@
Long-term time series forecasting (LTSF) (acmsurvey2, ; ACMtimerserissurvey, ) stands as a cornerstone within a multitude of practical applications for complex real-life scenarios, such as finance (DBLP:journals/ijon/TangSZYHJTL22financial, ), traffic, weather (bi2023accurate, ) and hospital admission rates prediction (DBLP:journals/midm/RamanADKCAAACIBM23, ), wielding the power to shape crucial outcomes. Witness the success of Transformer (vaswani2017attention, ) and further applications in numerous fields (li2022video, ; li2023dynamic, ; li2024contrastive, ), researchers have begun to utilize the Transformer and its evolutionary descendants (lee2024ts-fastformer, ; nie2022time, ; zhou2022fedformer, ; Zhou2020informer, ) as the de facto backbone in this intricate landscape. It showcases unparalleled prowess in handling extensive sequence data, elucidates patterns, and unveils latent structures that underscore sequential phenomena. More details, these Transformer-based LTSF systems have demonstrated strong performance, for which they can predict hundreds of time steps ahead, in contrast to other sequence processing networks and traditional methods that were mostly only capable of making predictions for 48-time points or less (Hochreiter1997LongSM, ; QinSCCJC17, ; wen2018multihorizon, ). Transformer-based architectures have demonstrated exceptional capability in modeling long-range dependencies, making them well-suited for LTSF (Wu2021autoformer, ).
Despite these advancements, it is important to recognize that prior LSTM-based models such as Long Short-Term Memory (LSTM)(Hochreiter1997LongSM, ) and its variants, including Seq2Seq(sutskever2014sequence, ) and Attention-based LSTM (bahdanau2014neural, ), approached LTSF differently by leveraging their innate ability to process sequential data. These methods excelled in capturing local temporal dependencies but struggled with learning long-range dependencies due to vanishing gradient issues. LSTM-based methods also relied heavily on carefully tuned window sizes for rolling forecasting, which inadvertently introduced redundancy similar to Transformer-based systems. However, unlike Transformers, LSTM methods lacked the architectural flexibility to mitigate the overfitting caused by redundant data, as increasing network depth often led to performance degradation rather than enhancement.
@@
Strikingly, when playing with these Transformer-based LTSF systems (lee2024ts-fastformer, ; Wu2021autoformer, ; Zhou2020informer, ), we observed that these models frequently attain their apex performance on validation datasets within a remarkably abbreviated span of training epochs, sometimes even the first epoch (Zhou2020informer, ), and then the performances dropped gradually. This observation hints at the latent underexploitation of the Transformer-based systems’ innate potential. The emergence of such a phenomenon can typically be attributed to a confluence of two primary factors: insufficient training data and the constrained parameter within the network architecture. However, the inherent modularity of these architectures allows for facile parameter augmentation by virtue of stacking additional Transformer modules. This architectural extensibility, coupled with the recognition that the constrained parameter space might not be the predominant bottleneck, guides our inference toward the more probable origin: the paucity of comprehensive and high-quality training data.
Existing models like Fedformer (zhou2022fedformer, ) and Informer (Zhou2020informer, ) have introduced specialized mechanisms to alleviate some limitations of redundancy in sequential data. Fedformer incorporates multi-scale decomposition to capture seasonal trends, while Informer uses a sparse self-attention mechanism to reduce computational overhead and focus on key temporal correlations. However, both approaches are still prone to the pitfalls of data redundancy stemming from rolling forecasting settings. Specifically, while sparse attention reduces the volume of redundant computations, it does not inherently address the similarity between consecutive training samples, which persists as an obstacle for longer sequences. Similarly, multi-scale decomposition focuses on improving representation but does not directly mitigate the negative impacts of overlapping inputs in rolling forecasting.
@@
Delving further into this phenomenon, we found that there is severe data redundancy within the training samples, stemming from the common practice of rolling forecasting settings (see Figure.1 for more details) during data augmentation. Under the rolling forecasting setting, if the stride size is 1, then the previous and current training samples are only one data point apart from each other. Since making longer predictions typically requires longer input sequences, training samples for LTSF models have a higher degree of similarity. For example, when the model’s input length is 2, the similarity between two consecutive training inputs is 50% (1/2); however, if we increase the model’s input length to 168, the similarity between two consecutive training inputs rises to 99.4% (167/168). Therefore, the LTSF models are more likely to suffer from overfitting due to such data redundancy. This issue significantly limits the diversity of training samples, reducing the models’ ability to generalize to unseen patterns, a challenge shared with LSTM-based models that rely heavily on overlapping inputs during training. One could argue that we could decrease the training sample similarities by increasing the stride size; however, the number of training samples decreases by the factor of the stride size, as the training sample numbers are calculated by:
(1) $N_{t}=\frac{N_{d}-S_{w}+1}{S_{s}}\rm{.}$
@@
Figure 1. Illustration of the rolling forecasting setting with stride size=1 and window size=6.
To address the aforementioned data redundancy issue, we propose CLMFormer, a novel pipeline with curriculum learning and a memory-driven decoder, designed to revitalize varying Transformer-based LTSF systems. Since data redundancy directly results from the heightened similarity present within the training data, our priority is to break this similarity while concurrently preserving the necessary training sample volume. To achieve this, we leverage the concept of curriculum learning, gradually increasing training difficulty and data variety by progressively introducing Bernoulli noise to the training data through a dynamic dropout scheme. Such a scheme reduces the propensity for overfitting while simultaneously enhancing the model’s capacity to capture sequential patterns. To further enhance the Transformer’s ability to recognize intricate sequential patterns from highly similar training data, we propose a memory-driven decoder to facilitate the forecasting process. This is accomplished by adopting a seasonal memory and memory-conditioned layer normalization operation, both of which effectively utilize past time-series information and introduce additional computations to enhance the model’s capability for complex pattern recognition. The main contributions of this paper are summarized as follows:
@@
The rest of this paper is organized as follows: Section 2 provides a comprehensive review of related work in LTSF and Transformer-based architectures. Section 3 presents the CLMFormer framework, detailing its progressive curriculum learning strategy and memory-driven decoder. Section 4 describes the experimental setup, including datasets, evaluation metrics, and baseline comparisons, followed by a thorough performance analysis. Section 5 discusses the limitations and potential future improvements of CLMFormer, while Section 6 concludes the paper by summarizing key findings and highlighting the broader impact of our approach.
2. Related Work
2.1. Time Series Forecasting
@@
Time series forecasting (TSF) (DBLP:journals/csur/BenidisRFWMTGBS23, ; TKDE3, ) is a critical task in various domains, encompassing domains such as finance, energy, healthcare, and more. Early approaches to TSF often revolved around traditional statistical and machine learning methods. These methods include ARIMA models (box1968somearime, ), exponential smoothing methods (gardner1985exponential, ), and seasonal decomposition of time series algorithms (cleveland1990stl, ). These techniques have been widely used due to their simplicity and interpretability. However, they often struggle to capture intricate temporal patterns and dependencies present in real-world data.
In recent years, deep learning techniques have gained traction for TSF. RNNs, including LSTM(Hochreiter1997LongSM, ), GRU(gru, ), and DeepAR(salinas2020deepar, ), have demonstrated their capability to capture temporal patterns. However, these models easily suffer from error accumulation, a phenomenon where small errors generated during earlier time steps can propagate and magnify as the sequence progresses(bengio2015scheduled, ). This issue can significantly impact the overall accuracy and stability of the model’s predictions, especially in long sequences or when dealing with complex relationships within the data. Despite these challenges, deep learning approaches have continued to evolve, showing remarkable advancements in multivariate time series forecasting (TKDECheng, ). Tang et al. (TangTKDE, ) introduced Enhanced-BF, a novel factorization approach for multivariate time series data, overcoming challenges like scattered data distribution through a three-phase process that improves robustness, efficiency, and accuracy in extracting meaningful features. Jin et al. (TKDEjin, ) addressed limitations of discrete neural architectures, computational complexity, and reliance on static graph structures by using dynamic graph neural ordinary differential equations to improve multivariant prediction performances. Beyond these developments, innovative architectures like TimeMixer (timemixer, ) and TimeMixer++ (wang2024timemixer++, ) have redefined time-series forecasting by leveraging multiscale-mixing and advanced pattern extraction techniques. TimeMixer disentangles seasonal and trend components across fine and coarse scales using Past-Decomposable-Mixing (PDM) and Future-Multipredictor-Mixing (FMM), achieving state-of-the-art performance in both short- and long-term forecasting. Building on this, TimeMixer++ introduces multi-resolution time imaging (MRTI) and multi-scale mixing (MCM) to extract comprehensive temporal and frequency-domain patterns, enabling superior adaptability across diverse tasks. However, these models require substantial data to train effectively, while real-world time-series datasets are often small, domain-specific, and strongly biased by trends, limiting their practicality and scalability in certain applications. Despite their sophistication, these models rely heavily on abundant and diverse training data. However, real-world time-series datasets often exhibit redundancy, particularly when rolling forecasting settings are applied, which undermines generalization and scalability.
2.2. Transformer-based LTSF Systems
Due to their superior performance in processing sequential data, Transformer-based models have achieved state-of-the-art results in numerous time-series tasks (wen2022transformers, ), particularly in long-term series forecasting (LTSF) problems (cao2023inparformer, ; Wu2021autoformer, ; Zhou2020informer, ; jin2023expressive, ). Research in this area has focused on several key directions. Some studies aim to reduce computational costs to enable longer predictions by employing sparse attention mechanisms, which reduce memory costs from O($L^{2}$) to as low as O($LlogL$), where L refers to the sequence length in attention layers (Wu2020AdversarialST, ; Zhou2020informer, ). Other works emphasize architectural improvements or enhanced patch representations to better learn seasonal and intricate temporal patterns.
Transformer-based models, however, still face challenges in sequential relationship representation and often suffer from slow training and inference speeds due to their deep encoder-decoder architectures. To address these issues, several innovative approaches have been proposed. For instance, Lee et al. (lee2024ts-fastformer, ) introduced the Sub Window Tokenizer, which simplifies and compresses input sequences with the help of a pre-trained encoder. The Autoformer model (Wu2021autoformer, ) introduced an Auto-Correlation mechanism, allowing sub-series level aggregation to effectively handle periodicity in time-series data. FEDformer (zhou2022fedformer, ), in contrast, adopts frequency-enhanced blocks and attention mechanisms, utilizing random subsets of frequency components along with learnable complex-number parameters to construct sparse temporal representations. InParformer (cao2023inparformer, ) integrates evolutionary seasonal-trend decomposition modules with an interactive parallel attention mechanism, combining frequency-aware and time-aware attention to better capture meaningful patterns in time-series data. PatchTST (nie2022time, ) improves representation learning by using subseries-level patches and promoting channel independence. Despite these advancements, Zeng et al. (zeng2023transformers, ) argued that Transformer-based LTSF models remain underutilized and proposed a simple linear model that surprisingly achieved superior performance. While these approaches effectively reduce model complexity and improve representation, they do not address the issue of training data redundancy caused by overlapping input sequences, particularly in rolling forecasting settings. Such redundancy limits the diversity of training samples and amplifies the risk of overfitting, which is especially problematic in long-term forecasting tasks.
The rise of deep Transformer networks, such as foundation models and large language models (LLMs), marks a significant paradigm shift in time-series forecasting. These models, originally developed for natural language processing tasks, have demonstrated unparalleled capacity for learning from massive datasets, capturing long-term dependencies, and adapting to diverse tasks. Foundation models, like the decoder-only model for time-series forecasting proposed by Das et al. (timesfoundationmodel, ), leverage their powerful pre-training capabilities to provide a unified architecture for forecasting across various applications. Similarly, Jin et al. (LLMintimeseriesforecasting, ) explored the use of existing LLMs to revolutionize time-series analysis, framing forecasting tasks as sequence generation problems and benefiting from the transfer learning capabilities of these large-scale models.
Despite their promise, employing these large models in time-series forecasting presents significant challenges. Unlike natural language tasks, time-series datasets are often small, domain-specific, and strongly influenced by underlying trends and seasonality. These characteristics make time-series data inherently biased, limiting the generalization capabilities of large models pre-trained on diverse datasets. Moreover, the massive scale of foundation models and LLMs demands extensive computational resources and vast quantities of training data, which are often unavailable in time-series domains. This data scarcity, combined with the unique and predictable nature of time-series patterns, makes training such large Transformer networks difficult and inefficient. Consequently, while these models hold great potential, their practical application in time-series forecasting must address the trade-offs between model complexity, data requirements, and the specialized characteristics of time-series data. In this paper, we revisit these models and uncover the issue attributed to data redundancy stemming from rolling forecasting settings. Most importantly, data redundancy caused by rolling forecasting settings persists as a key bottleneck in long-term forecasting systems, compounding challenges associated with small, domain-specific datasets. This redundancy not only exacerbates overfitting but also diminishes the representational quality of training data, creating a significant gap in existing Transformer-based LTSF solutions. To bridge this gap, CLMFormer directly tackles data redundancy by employing curriculum learning to diversify training samples and a memory-driven decoder to enhance seasonal and dependency modeling, ensuring robust performance on long-term time-series forecasting tasks.
3. Methodology
@@
In this section, we present our solutions to revitalize transformer-based systems that specifically target the challenges faced by LTSF. Our approach is based upon the Transformer architecture; we aim to enhance Transformer models on LTSF problems by making improvements to mitigate data redundancy from both training optimization prospects and the architecture. The overview of our approach is illustrated in Figure.2. Our model comprises three major elements, namely the input embedding, the main Transformer-based model, and the seasonal memory component, whereby we propose to train the model based on a progressive training strategy. We first introduce the background of employing Transformer models for LTSF and then describe the training strategy. The seasonal memory component is described in the remaining subsections.
@@
Typical Transformer-based LTSF models normally train under rolling forecasting settings with a fixed window and a stride size = 1, as shown in Figure.1. The input data at time t can be written as $X^{t}=\{x_{1}^{t},x_{2}^{t}...x^{t}_{\emph{L}_{S}}|x_{k}^{t}\in\rm{R}^{\emph{d}_{\emph{f}}}\}$, where ${L}_{S}$ is the input length and the feature dimension $d_{f}>1$. Moreover, the model output is $Y^{t}=\{y_{1}^{t},y_{2}^{t}...y^{t}_{\emph{L}_{p}}\}$, where ${L}_{p}$ is the prediction length, and the output feature dimension is flexible and is not limited to single sequence predictions. Algorithm.1 presents the training process of the whole system.
Algorithm 1 Training Process of Our CLMformer
0:  Time-series input data $X$, maximum dropout rate $\theta_{max}$, learning rate $\alpha$, hyperparameter $\gamma$
0:  Trained model parameters $\Theta$
1:  Initialize: Memory matrix $M_{0}$, model parameters $\Theta$
2:  for epoch = 1 to $N_{\text{epochs}}$ do
3:     for batch $B$ in training set do
4:        Extract input sequence $X_{B}$ and target sequence $Y_{B}$
5:        Curriculum Learning Strategy:
6:        Compute dynamic dropout rate
7:        Apply dropout to input sequence: $X_{B}\leftarrow\text{Dropout}(X_{B},\theta_{d}(t))$
8:        Encoder Processing:
9:        Encode $X_{B}$ using Transformer encoder to obtain hidden representation $H$
10:        Memory-Driven Decoder Processing:
11:        for time step $t$ in decoder do
12:           Compute attention-weighted memory update
13:           Apply Memory-Driven Conditional Layer Normalization
14:           Compute final decoder output $\hat{Y}_{t}$
15:        end for
16:        Compute loss $\mathcal{L}(Y_{B},\hat{Y}_{B})$
17:        Update model parameters: $\Theta\leftarrow\Theta-\alpha\nabla_{\Theta}\mathcal{L}$
18:     end for
19:  end for
20:  Return trained model parameters $\Theta$
3.1. Base Model
3.1.1. Input Embedding